OpenStudy (anonymous):

Multiply and simplify if possible. sqrt6 * sqrt2 I tried to solve it the way it looked and was way off.

OpenStudy (solomonzelman):

When you multiply square roots. \[\rm RULE:~~~\bf \sqrt{x} \times \sqrt{y}=\sqrt{x \times y}\]\[\bf \sqrt{6} \times \sqrt{2}=?\]

OpenStudy (anonymous):

so it would be \[2\sqrt{3}\] ?

OpenStudy (solomonzelman):

AWESOME, great job! And you took the 4 out of the root ... !!!!!!!!!

OpenStudy (solomonzelman):

You seem to actually now this stuff ;)

OpenStudy (anonymous):

ahhhh satisfaction! lol. thank you! :) wat do you mean? lol.

OpenStudy (anonymous):

what*